OpenStudy (anonymous):

Find the value of m so that Integral from m to infinity e^(-x/2)dx = 1/2

OpenStudy (anonymous):

\[\int\limits_{m}^{\infty} e^(-x/2) dx =1/2\]

OpenStudy (anonymous):

i guess the anti derivative is \(-2e^{-\frac{x}{2}}\)

OpenStudy (anonymous):

as x goes to infinity this goes to zero right?

OpenStudy (anonymous):

so your real job is to set \[\frac{1}{2}e^{-\frac{x}{2}}=\frac{1}{2}\] and solve for \(x\)

OpenStudy (anonymous):

make sense? and is it clear what to do now?

OpenStudy (anonymous):

How did you get the 1/2 e^-x/2

OpenStudy (anonymous):

the first time or the second time?

OpenStudy (anonymous):

second time

OpenStudy (anonymous):

first you want \[\int e^{-\frac{x}{2}}dx\] the anti - derivative. i get by a mental u - sub \[-\frac{1}{2}e^{-\frac{x}{2}}\] so far so good?

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

then to compute \[\int_ m^{\infty}e^{-\frac{x}{2}}dx\] you take the anti derivative you compute \[\lim_{t\to \infty}-\frac{1}{2}e^{-\frac{x}{2}}-\left (-\frac{1}{2}e^{\frac{m}{2}}\right)\]

OpenStudy (anonymous):

now the limit is part is clearly zero because \(e^{-\frac{x}{2}}\) goes to zero lickety split

OpenStudy (anonymous):

so you are left with \[\frac{1}{2}e^{-\frac{m}{2}}\] and you want that to be \(\frac{1}{2}\) so write \[\frac{1}{2}e^{-\frac{m}{2}}=\frac{1}{2}\] and solve for \(m\)

OpenStudy (anonymous):

k?

OpenStudy (anonymous):

Sweet thanks so much!

OpenStudy (anonymous):

yw